### 12 Dec 2023 v1

# On Diverse Preferences for Large Language Model Alignment

Abstract
The alignment of large language models (LLMs) with human values is crucial for the development of artificial general intelligence (AGI). One promising approach to achieve this alignment is reinforcement learning from human feedback, which employs a reward model (RM) learned from human preference datasets to guide LLMs in generating text that aligns with human preferences. Through intensive experiments and analysis of reward distribution, this paper finds that preference datasets are diverse from each other, even though they are all proposed to align human preference. Hence, mixing diverse human preference datasets to increase data size for enhancing reward modeling could fail. To address the issue and capture the shared human values from diverse preferences, a new training policy called MORE is introduced, which minimizes preference bias by adaptively adjusting the preference objective across diverse preferences. Experiments with the Pythia-1.4B model and five mixed preference datasets show that MORE achieves superior reward accuracy and lower calibration error, highlighting its ability to leverage diverse human preference data.
Machine Learning, ICML
1 Introduction
[…]
The emergence of large language models (LLMs) such as ChatGPT and LLama series (Touvron et al., 2023a, b) have propelled us a step forward toward achieving artificial general intelligence (AGI). One key factor in the success of large models is aligning them with human preferences (Christiano et al., 2017), enabling these models to better meet societal expectations, ensuring that their outputs are more useful, benign, and aligned with shared values. Among the alignment algorithms, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has gained prominence as an approach to address the challenge of learning from human preferences. The training process aims to maximize the reward provided by a reward model (RM), which is learned from paired preferred and dispreferred data samples (Bai et al., 2022c; Ganguli et al., 2022; Fernandes et al., 2023). Alternative algorithms to PPO have been continuously proposed (e.g., RLAIF (Lee et al., 2023), ReST (Gulcehre et al., 2023), Rrhf (Yuan et al., 2023), DPO (Rafailov et al., 2023), APO (Cheng et al., 2023b), etc.), bringing to the emergence of numerous outstanding deployable LLMs, such as Anthropic’s Claude (Bai et al., 2022b), Microsoft’s Bing Chat, and Google’s Bard (Manyika, 2023).
The effectiveness of these RLHF series algorithms depends on high-quality RMs (Ouyang et al., 2022). The training of RMs requires a large number of human preference data collected from the real world (typically from crowdworkers) (Bai et al., 2022b). Such human preference data generally needs to be helpful, harmless, benevolent, and unbiased. However, in a pluralistic world, individuals can have different preferences toward the same response due to their diverse experiences, educational backgrounds, religious beliefs, and social values. Moreover, even for the same individual, the value assigned to a specific response from a language model can vary depending on the application scenario (Cheng et al., 2023a). Hence, data annotated by individuals with diverse preferences inherently encompasses a range of opinions and biases. Therefore, leveraging this preference data effectively to train a robust RM poses a significant challenge (Cheng et al., 2023a). The issue arising from these preference divergences or conflicts is of paramount importance, yet it has received little attention and dedicated research thus far in reward modeling.
To identify the challenge, we first investigate a broader range of preferences to understand the inherent characteristics of preference data. We conduct a series of experiments using five open preference datasets, including Helpful&Harmless (Bai et al., 2022a), Webgpt (Nakano et al., 2021a), OASST1 (Köpf et al., 2023), and Summarize (Stiennon et al., 2020). Taking each dataset as a preference with characteristics, we solely train a set of RMs marked by $\{\theta_{\text{helpful}},\theta_{\text{harmless}},\theta_{\text{oasst}},\theta_{\text{webgpt}},\theta_{\text{summ.}}\}$ on each corresponding datasets and testing them on same preference samples as shown in Figure 1. We observe:
  • The implicit preferences across datasets are significantly different. Although these datasets are presented for LLMs to align with human values, the RMs provide very diverse reward values on the same preference data samples. The phenomenon reflects the preference inconsistency hidden in the datasets.
In this paper, we conceptualize the output of RM as a shared reward, accompanied by an additional reward drift. The shared reward indicates the shared human preferences and the reward drift indicates the characteristics. To capture the shared reward in reward modeling, we introduce a straightforward training policy called MORE. The MORE policy aims to train an RM across multiple comparison datasets, enabling the model to capture a broader range of preferences and reduce the impact of reward drift. Our main contributions are encapsulated as follows:
  • We conduct a thorough analysis of how diverse preference datasets affect reward modeling, focusing on the performance of the reward model and the distribution of rewards. Our findings indicate that diverse preference datasets lead to significant variations in reward distributions within the models. Consequently, these variations negatively impact the performance of the reward model.
  • We frame the reward as a shared reward plus reward drift. We propose a straightforward training policy called MORE for reward modeling. This policy only modifies the data sampling and loss definition, adaptively updating the RM by minimizing reward drift.
  • We conduct experiments on the Pythia-1.4B model (Biderman et al., 2023) to evaluate the effects of reward modeling on diverse datasets. Through these experiments, RM trained by MORE achieves the best reward accuracy and the lowest expected calibration error value (Zhu et al., 2023), compared with hybrid and sequential training policies. The results indicate it can output more accurate reward values for the RLHF task.
Figure 2: Overview of MORE policy. The MORE policy consists of the following steps: 1) It collects a diverse batch of data from the mixed dataset. 2) It calculates the gradient about the weights of the reward head. 3) It solves a minimization problem to determine the scalarization factor $\lambda$ for MORE loss. 4) Finally, update the RM via utilized optimizer.
2 Preference Modeling on Diverse Preference Datasets
We study the reward modeling task on diverse preference datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$. The $K$ denotes the number of diverse preferences and each preference is represented by a set of preference data samples $\mathcal{D}_{i},i\in[K]$. The problem setting covers the case that a single preference dataset implicitly preserves different preferences, such as Helpful&Harmless (Bai et al., 2022a). It also covers the case that mixed different preference datasets were collected in various tasks (Dong et al., 2023), domains (Cheng et al., 2023a), or individuals (Bai et al., 2022b).
2.1 Preliminaries
Preference alignment pipeline. A RM parameterized by $\theta$ can be denoted as a mapping $r_{\theta}:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, which provides a real-valued reward score $r_{\theta}(x,y_{w})$ qualifying a textual response $y=(y_{1},y_{2},\dots,y_{M})\in\mathcal{Y}$ corresponding to an input prompt $x=(x_{1},x_{2},\dots,x_{N})\in\mathcal{X}$. Given a sample $(x,y_{w},y_{l})\sim\mathcal{D}$ from preference dataset $\mathcal{D}$, $r_{\theta}$ is expected to provide a preference score with $r(x,y_{w})>r(x,y_{l})$, indicating the response $y_{w}$ is preferred. The Bradley-Terry model (David, 1963) represents the preference function $p(y_{w}\succ y_{l}|x)$:
$p(y_{w}\succ y_{l}|x)=\frac{\exp{r(x,y_{l})}}{\exp{r(x,y_{l})}+\exp{r(x,y_{w})}},$
where $\sigma(\cdot)$ is an activation usually set as the Sigmoid function. The standard reward modeling process optimizes a pre-trained language model on preference dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{n}$ with the following binary rank loss,
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\log(\sigma\left(r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})\right)).$ (1)
Using the reward $r_{\theta}(x,y)$ in the RLHF objective is to optimize for the policy $\pi$ that maximizes the expected reward while minimizing the distance between $\pi$ and some reference policy $\pi_{\text{ref}}$ as:
$\max_{\pi}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi(y|x)}\left[r_{\theta}(x,y)-\beta\mathbb{D}_{\text{KL}}[\pi(y|x)\|\pi_{\text{ref}}(y|x)]\right].$ (2)
2.2 Preference Bias as Reward Drift.
We use $r^{*}(\cdot,\cdot)$ to denote the shared preference reward function, which always gives proper reward value to the contents with the shared values across human beings. As the collected human-feedback datasets are limited and implicitly biased, they usually can not cover the whole shared value. Therefore, training an RM $r_{\theta}$ on a limited preference dataset can be viewed as drifting from an optimal reward. We form the reward drift of a reward model parameterized by $\theta$ from a data level:
$r_{\theta}(x,y)=r_{\theta}^{*}(x,y)+\tilde{r}_{\theta}(x,y),\;\forall(x,y)\in\mathcal{X}\times\mathcal{Y},$ (3)
where $\tilde{r}_{\theta}(x,y)$ is the reward drift learned from the given dataset. Then, we investigate the vanilla ranking loss for reward modeling. Substituting reward function in equation 1 with the drifted form equation 3, we have
$\mathcal{L}_{\text{rank}}(\theta;\mathcal{D})$ $=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}[\log(\sigma(r_{\theta}^{*}(x,y_{w})-r_{\theta}^{*}(x,y_{l})$ (4)
$\quad+\tilde{r}_{\theta}(x,y_{w})-\tilde{r}_{\theta}(x,y_{l})))].$
Hence, updating the RM to minimize the rank loss will enlarge the reward differences (input of the Sigmoid function). Simultaneously, the reward drift is also enlarged.
2.3 Reward Drift Modeling
We discuss the unique form of learning general preferences from diverse datasets in this part. Firstly, we justify the reward agreement and disagreement in preference modeling:
Definition 1 (Reward agreement and disagreement).
Given a set of diverse preference datasets $\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$, we can obtain RM sets $\{\theta_{1},\dots,\theta_{K}\}$, which are trained solely on every single dataset. Then, given an comparison sample $(x,y_{w},y_{l})\in\mathcal{X}\times\mathcal{Y}$, we say this is an agreed sample if $r_{\theta_{j}}(x,y_{w})>r_{\theta_{j}}(x,y_{l})$ for all $j\in[K]$. Otherwise, if there is a pair $(i,j)$ such that $r_{\theta_{i}}(x,y_{w})>r_{\theta_{i}}(x,y_{l})$ and $r_{\theta_{j}}(x,y_{w})<r_{\theta_{j}}(x,y_{l})$, this sample is disagreed.
Compared with the previous discussion (Davani et al., 2022; Basile et al., 2021; Wang et al., 2023a; Zhao et al., 2023a) in label disagreement, we are focusing on the rewards disagreement across the RMs.
Letting $\theta$ be the RM trained on mixed diverse datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$, the $r_{\theta}(x,y)$ can be viewed as a multi-task learner with shared parameters (Sener & Koltun, 2018). Then, the reward value provided by $r_{\theta}(x,y)$ can be decomposed into voting format weighted by an implicit $\lambda$:
$r_{\theta}(x,y)=$ $\sum_{i=1}^{K}\lambda_{i}r_{\theta_{i}}(x,y)$ (5)
$=$ $r_{\theta}^{*}(x,y)+\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y),\forall(x,y)\in\mathcal{D},$
where the shared reward $r^{*}_{\theta}(\cdot,\cdot)$ is the same with arbitrary $\lambda$, and $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is the reward drift. We interpret that the $\tilde{r}_{\theta_{i}}(\cdot,\cdot)$ is provided by subset of parameters $\theta_{i}$, representing the preferences from the $i$-th dataset $\mathcal{D}_{i}$. For example, the $\theta$ can be implemented as an ensemble model, where $\{{\theta_{i}}\},i\in[N]$ is the base models. Therefore, it is natural to adjust the weight $\lambda$ in an ensemble manner (Coste et al., 2023; Jang et al., 2023; Touvron et al., 2023a) to mitigate the reward drift such that $\min\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)$. Moreover, if the datasets are extremely diverse, the data samples are very likely to be disagreed by $K$ diverse preference by Definition 1. Hence, there always exists an combination $\lambda$ such that $\sum_{i=1}^{K}\lambda_{i}\tilde{r}_{\theta_{i}}(x,y)=0$.
2.4 Training Policy: MORE
Compared with an ensemble of a set of RMs (Jang et al., 2023), we are more interested in training a single RM that learns the shared preference. Our idea is to reduce the model update on reward drift during RM training via linear scalarization (Barrett & Narayanan, 2008).
MORE loss function. Inspired by previous analysis, finding proper weights $\lambda$ for mitigating reward drifts is a feasible way. Analogously, we propose training an RM to capture the shared preference distribution across multiple datasets with the training objective:
$\mathcal{L}_{\text{MORE}}(\theta;\mathcal{D})=\sum_{i=1}^{K}\lambda_{i}\mathcal{L}_{\text{rank}}(\theta,\mathcal{D}_{i}),$ (6)
where $\sum_{i}^{K}\lambda_{i}=1$. Compared with vanilla ranking loss in Equation 1, the above loss additionally focuses on the combination relation across preferences. It suggests training a better RM via adjusting training weights $\lambda$. Moreover, this formulation also covers typical training strategies. For example, directly mixing diverse preference datasets $\mathcal{D}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$ and training a RM induces $\lambda_{i}=\frac{|\mathcal{D}_{i}|}{|\mathcal{D}|}$ (McMahan et al., 2017). Therefore, if the number of data samples from a single preference is greatly larger than other preferences, the RM is likely to drift to the preference with more samples. Excluding data quantity, the weight is also decided by the quality of data samples in the training process (Katharopoulos & Fleuret, 2018).
Computing $\lambda$. We propose to use adaptive weighting methods to balance the bias across different domains and adjust the reward modeling process. Straightforwardly, the mitigation task in Equation 9 can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener & Koltun, 2018; Zhou et al., 2022; Zeng et al., 2023), while the computing cost for solving it is proportional to the size of parameters $\theta$. To avoid expensive computation, we only utilize gradients of the reward head $\theta_{\text{rm}}\in\mathcal{R}^{h}$ from each preference. In detail, we obtain the hidden states $z=r_{\theta_{\text{lm}}}(x)$ before the reward head and compute the gradient of the reward head solely with data $(z,y_{w},y_{l})$. In summary, the $\lambda$ is computed by:
$\lambda=\arg\min_{\lambda}\left\|\sum_{i=1}^{K}\lambda_{i}g_{i}\right\|^{2},s.t.,\;\sum_{i=1}^{K}\lambda_{i}=1,\lambda_{i}\geq 0,$ (7)
where $g_{i}$ is a momentum gradient related to $\nabla\mathcal{L}_{\text{rank}}(\theta_{\text{rm}})$. During the training process $t\in[T]$, the gradients is marked with $\{g_{1}^{t},\dots,g_{K}^{t}\},t\in[T]$. Moreover, gradients of the reward head are cached and updated in a momentum manner with $\alpha$:
$g_{i}^{t}=(1-\alpha)g_{i}^{t-1}+\alpha\frac{1}{|\mathcal{B}_{i}|}\sum_{\mathcal{B}_{i}}\nabla_{\theta_{\text{rm}}}\mathcal{L}_{\text{rank}}(\theta_{\text{rm}};\mathcal{B}_{i}).$ (8)
The cached gradient $g_{i}^{t}$ preserves the history updating information. It stabilizes the results $\lambda^{t}$ by involving the information from the past data samples (Zeng et al., 2023).
Outline. The MORE loss requires simple modification on batch data sampling and additional computing $\lambda$. In detail, the pipeline of MORE is given in Figure 2 and described as follows: 1) Sample a diverse batch data =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{K}$, $\mathcal{B}_{i}=\{x,y_{w},y_{l}\}_{b=1}^{B}$ and input the batch data forward the RM and obtain the hidden states $\{z_{i}\}_{b=1}^{B}$ before the reward head $\theta_{\text{rm}}$. 2) Compute the gradient of reward with data $\{z,y_{w},y_{l}\}$ and update the cache $\{g_{i}^{t}\}$ following equation 8. 3) Compute the weights $\lambda$ by equation 7. Finally, we substitute the loss weights in equation 6 as the final loss for the optimizer to conduct backward and model updating. This procedure mitigates the potential reward drift from the overall reward.
Theoretical explanation The MORE loss minimizes the ranking loss after solving a reward drift mitigation task. Let batch data =$\mathcal{B}=\{x^{(b)},y_{w}^{(b)},y_{l}^{(b)}\}_{b=1}^{B}\sim\mathcal{D}$ be the sampled batch data from diverse datasets. Furthermore, $\mathcal{B}_{i}\sim\mathcal{D}_{i}\subset\mathcal{B},\forall i\in[K]$ is the subset of batch data from the $i$-th preference dataset. We have the gradient of MORE loss:
$\nabla_{\theta}\mathcal{L}_{\text{MORE}}(\theta;\mathcal{B})=\sum_{b=1}^{B}\left[-\nabla_{\theta}\log(\sigma(r_{\theta}^{*}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)})))\right]$ (9)
$\quad+K\underbrace{\min\sum_{i=1}^{K}\lambda_{i}\sum_{j=1}^{|\mathcal{B}_{i}|}\left[-\nabla_{\theta}\log(\sigma(\tilde{r}_{\theta}(x^{(j)},y_{w}^{(j)}-y_{l}^{(j)})))\right]}_{\text{Reward Drift Mitigation Task}},$
where we let $r_{\theta}(x,y_{w}-y_{l})=r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$ and adjust $\lambda$ (equation 7) to minimize the effect of reward drifts. Furthermore, $\mathcal{L}_{\text{MORE}}$ shares the same magnitude of vanilla loss function $\mathcal{L}_{\text{rank}}$ in expectation over the whole training data set. The point can be justified by comparing it with vanilla ranking loss as shown in Appendix A. The main difference falls in the loss of the mini-batch descending process.
3 Experiments
In this section, we conduct an empirical evaluation of MORE for training RMs on diverse preference datasets. Firstly, we provide a detailed description of our experimental setup in Section 3.1. Then, we demonstrate the performance enhancement achieved by employing MORE in contrast to traditional RM training policies in Section 3.2. Next, we examine the distribution of reward values generated by the trained RMs. Through analysis of reward values, we demonstrate how the diverse preference dataset affects reward modeling.
Table 1: Reward model performance on diverse datasets.
Training Testing Dataset (Acc %) Metrics
Dataset Policy Helpful Harmless Oasst1 Webgpt Summ. Avg. Acc W.Avg. Acc $*{}^{*}$
- Raw 49.78 47.18 51.15 49.84 49.88 49.56 49.47 0.239
Helpful - 66.75 35.89 62.85 60.06 53.34 55.77 55.68 0.179
Harmless - 35.67 67.21 43.10 38.41 47.18 46.31 47.06 0.296
Oasst1 - 59.09 43.29 66.61 66.31 55.82 58.22 57.47 0.119
Webgpt - 60.96 41.13 64.00 67.37 49.11 56.51 55.32 0.128
Summ. - 60.07 42.56 58.68 60.36 58.44 56.02 56.18 0.206
ALL Ensemble 60.92 44.89 65.83 67.83 56.82 59.26 56.88 0.139
ALL Hybrid 57.14 52.63 66.30 66.15 58.25 60.09 59.60 0.078
ALL Sequential 57.95 53.71 66.71 67.68 57.07 60.62 60.36 0.084
ALL MORE 58.46 61.85 65.88 66.31 59.28 62.35 61.68 0.074
  • a
    $*{}^{*}$ower is better.
3.1 Experiment Setup
Datasets. We utilize open human preference datasets for reward modeling, which include Helpful&Harmless (Bai et al., 2022a), OASST1 (Köpf et al., 2023), Webgpt (Nakano et al., 2021a), and Summarize (Stiennon et al., 2020). To avoid the preference unbalance from the size of training sets, we randomly sample 17,106 samples from Helpful, Harmless, OASST1 and Summarize training sets to align with the Webgpt dataset. Besides, we note the unbalance of training data can be addressed by resampling methods (Estabrooks et al., 2004). The corresponding datasets contain (2,352, 2,312, 957, 901, 3,620) test samples respectively.
Baselines. We compare our method with conventional fine-tuning strategies (Dong et al., 2023) in the training language models, which include sequential training and hybrid training. We note that hybrid training policy is essentially a special case of MORE with a static $\lambda=\frac{1}{K}$. Moreover, the sequential training policy only samples a batch of data from single preference datasets during training. We also compare with the naive ensemble, which merely averages the reward values of independently trained rewards following equation 5.
Evaluation metric. We use the preference accuracy on test datasets for each domain. If an RM $r(x,y_{w})>r(x,y_{l})$ for a test sample $(x,y_{w},y_{l})$, we denote it as a correct prediction. The preference accuracy is then computed as the proportion of correct predictions within all testing response pairs. However, preference accuracy only provides pairwise comparisons of responses and does not reflect the degree of preference for each response. Following Bai et al. (2022b); Cheng et al. (2023b), we examine the probability calibration to test if the learned RMs accurately represent the human preference distribution. This is measured by the Expected Calibration Error (ECE) (Naeini et al., 2015; Zhu et al., 2023).
Training setup. Due to the limitation of computational resources, we only select Pythia-1.4B (Biderman et al., 2023) as the LM base for RM training. Experiments with a large scale of the base model are in progress. We use the last token embedding of the output hidden states as the pooled hidden representation, then add one linear layer (RM head) with the scale-value output on it to predict reward scores. All RM training batch size is set to 64. The max input sequence length is 512. All RMs are fine-tuned with one epoch. We use optimizer AdamW with learning rate $5e^{-5}$. We set $\alpha=0.5$ for MORE.
Table 2: Sequence of Harmless dataset potentially affects the performance. Seq.@$i$ indicates the Harmless is $i$-th utilized dataset in sequential training.
Training Testing Dataset (Acc %) Metrics
Dataset Policy Helpful Harmless Oasst1 Webgpt Summ. Avg. Acc W.Avg. Acc ECE
ALL Seq.@1 56.50 59.64 65.88 65.54 59.22 61.35 61.06 0.107
ALL Seq.@2 58.80 59.29 66.40 67.22 58.01 61.94 61.64 0.075
ALL Seq.@3 58.24 53.20 64.31 67.22 58.34 60.26 60.00 0.087
ALL Seq.@4 57.95 53.71 66.71 67.68 57.07 60.62 60.36 0.084
ALL Seq.@5 60.45 60.51 67.03 66.31 58.06 62.46 62.20 0.076
ALL MORE 58.46 61.85 65.88 66.31 59.28 62.35 61.68 0.074
Figure 3: Statistics of reward values on testsets provided by the RM trained with different policies on all mixed datasets. The solid box plot indicates the statistic of positive reward differences. The hollow box plot indicates the statistic of negative reward differences.
3.2 Reward Modeling on Multiple Preference Datasets
This section evaluates the performance of the RM trained on diverse preference datasets. The results are given in Table 1. Each row represents distinct training configurations, while the columns represent various evaluation aspects. The term “Avg.” denotes the arithmetic mean of accuracy across all test domains, while “W.Avg.” signifies the weighted average of accuracy, considering the size of the respective testsets.
Main performance. The MORE achieves better testing accuracy and lower ECE values, indicating that the policy provides an accurate reward value on the test. In the last four rows of the table, we train an RM on mixed datasets with different policies. The ensemble, hybrid, and sequential policies show similar test accuracy on single domains and average accuracy. However, the Harmless dataset, which has different preferences compared to the other four datasets, experiences the largest accuracy degeneration with these policies. In contrast, MORE exhibits the lowest performance degeneration on the Harmless test.
Agreement & disagreement analysis. The agreement or disagreement of preferences can be reflected in the testing accuracy. These findings highlight the importance of identifying and understanding the agreement or disagreement across collected datasets in practical applications. By comparing the experimental results of the first six rows, we can observe that when the model is trained on datasets from other domains, the test accuracy on the Harmless dataset decreases. This indicates a disagreement in preferences between the Harmless dataset and the other datasets.
In contrast, we observe a phenomenon of agreement when training the RM on one domain, which leads to an increase in test accuracy on the other preferences. For example, training the RM on any of the preference datasets such as Helpful, Oasst1, Webgpt, or Summarize will increase the test accuracy on the other preferences, even if the model has never seen the preference data samples.
Sequential training analysis. We also find that sequential training is sensitive to the sequence of preference datasets, as shown in Table 2. When adjusting the position of the Harmless dataset while maintaining a fixed sequence for the other datasets (Helpful, Oasst1, Webgpt, Summarize), we observe that sequential training slightly outperforms MORE in terms of testing accuracy when the Harmless dataset is placed at the end. However, MORE performs better on the ECE metric. Therefore, while sequential training has the potential to achieve comparable performance with MORE, it requires additional tuning efforts to adjust the dataset sequence, which can be troublesome when dealing with large and diverse datasets.
Why sequential training works better when Harmless at the last? The training performance is influenced by the proportion of agreement and disagreement data. Previous literature (Touvron et al., 2023a; Bai et al., 2022c) has shown that there is a disagreement between the Helpful and Harmless datasets. Additionally, our observations indicate that the Oasst1, Webgpt, and Summarize datasets share distributional similarities with Helpful (Figure 1) and exhibit similar testing accuracy (Table 1). As a result, the Harmless dataset disagrees with the other datasets, and its proportion in the mixed datasets is relatively small. Due to the smaller proportion of Harmless data, the model updating process allocates less attention to this dataset compared to the others. Sequential training can achieve better performance by adjusting the sequence of datasets. For instance, the recently seen samples are less likely to be forgotten (Wang et al., 2023b). In contrast, MORE successfully enhances performance by adaptively re-weighting losses of different preferences during training using Equation 6. MORE automatically captures the agreement and disagreement through Equation 9 and balances the model updating process accordingly.
3.3 Analysis on Diverse Distribution of Reward Values
Reward accuracy and ECE only measure the quality of reward values provided by RMs. How preference diversity plays into the reward model has not been discussed. This section presents our analysis on the statistics of reward values, highlighting the impact of diverse preferences on reward models.
Experiment setup. We first partition the testset into an “agreement testset” and a “disagreement testset”. In detail, we let $\{\theta_{\text{helpful}},\theta_{\text{harmless}},\theta_{\text{oasst}},\theta_{\text{webgpt}},\theta_{\text{summ.}}\}$ represent a preference model set. We utilize each preference model in the set to infer every data sample from the testset. We obtain the agreement testset which comprises data samples that agreed by the preference model set following Definition 1. The remaining samples form the disagreement testset. We analyze statistics and distribution of the reward difference (i.e., $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$) in reward values. It is worth noting that a positive reward difference $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})>0$ indicates that the model accurately assigns a higher reward value to the winning sample.
Analysis of reward differences on each preference. Our findings indicate that the reward value distributions of different reward models are generally distinct. When examining the RM trained on different single preference datasets, we observe a significant diversity in reward values. In Figure LABEL:fig:reward_dist, we present the reward difference values of the first 50 test samples from the Harmless dataset as an example (Figure LABEL:fig:reward_a). Notably, even on the agreed samples, the baseline model provides different rewards. This observation suggests that the preference distribution learned from a single dataset by the reward model is highly diverse compared to others. Furthermore, Figure LABEL:fig:reward_b provides statistics of the reward difference across the entire test set, further supporting our statements regarding the diversity of reward values.
In brief conclusion, despite presenting all the datasets to align with human preferences, the reward models trained on each dataset produce diverse results for the same data samples. This phenomenon suggests that the human preferences captured in these datasets are diverse from each other. Additionally, the difficulty of training the reward model on different preference datasets varies. Furthermore, even when a reward model correctly assigns rewards to a pair of comparison samples, the reward values can still exhibit significant differences. These observations highlight the need for further research on the task of reward modeling and the distribution of reward values. Such studies may help improve the accuracy of reward models in accurately capturing the degree of preferences.
Analysis of reward difference on training policies. MORE prevents the RM from outputting extreme reward values and maintains the reward accuracy (shown in Table 1). Specifically, we observe the reward values provided by RMs trained with different policies, as depicted in Figure 3. Compared with the sequential training policy, MORE outputs reward values and variance with lower mean and variance. Connecting with our methodology, MORE mitigated the reward drifting. Hence, it tends to output a lower absolute reward signal as more accurate reward values according to ECE in Table 1. Furthermore, the hybrid training policy also yields lower reward values compared to MORE. However, the reward accuracy is poorer in this case, which indicates that compressing the mean and variance of reward differences may degenerate the performance. These results suggest that the too-high or low reward differences may not be good in testing metrics. In contrast, MORE generates appropriate reward values, thereby enhancing the quality of reward modeling.
4 Discussion
We have proposed a simple yet effective method for RM to enhance the performance of learned RMs using diverse preference datasets. Our evaluation focuses on assessing the capabilities of trained RMs without completing the RLHF pipeline, which leaves us with some unanswered questions for future work.
Limitations. In this version, we have exclusively evaluated the MORE approach on the base model Pythia-1.4B. Further works on larger or smaller models are in progress. Besides, we randomly sample training samples from part of the dataset to avoid the effects of unbalanced datasets. The experiment of MORE handling the unbalancing mixed datasets will be presented in the next version. Throughout our experiments, we have found it necessary to utilize equation 8 in MORE, where we set $\alpha=0.5$ for all experiments. Further sensitive study on the value of setting $\alpha$ is in progress.
Future work. We have demonstrated the alignment ability of the reward model using the ECE metric (Zhu et al., 2023). However, the ultimate performance of the LLM, fine-tuned by the RM, remains a crucial aspect of our investigation. In our future work, we plan to complete the RL component in the RLHF pipeline. Additionally, we aim to address several key questions regarding the reward model and MORE: 1) How do the number of diverse datasets and the scaling of base models impact the performance of MORE? 2) Does the RM trained by MORE yield improved results in RL tasks? 3) Can learning feedback from mixed diverse preference datasets enhance the performance of the LLM? Furthermore, we will delve deeper into understanding the implications of preference diversity in reward modeling. Gaining insights into the effects of preference diversity is crucial for designing algorithms that exhibit enhanced data efficiency.
Extension and Applications. The concept of learning shared preferences from mixed diverse preference datasets can be extended to DPO by incorporating the MORE loss (Rafailov et al., 2023). Furthermore, the approach of learning shared preferences through MORE can be implemented as a preference modeling pre-trained (PMP) paradigm (Askell et al., 2021). Subsequently, the RM trained using MORE can be easily fine-tuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications and opens up new possibilities for preference modeling.
5 Related Work
RLHF framework has become the most popular techniques (Leike et al., 2018; Nakano et al., 2021b; Ouyang et al., 2022; Bai et al., 2022b) to align language models towards helpfulness, and harmlessness. They all utilize an RM to align machine learning systems with human performance, which directly decides the performance of preference alignment. As the RM is the most important component in the RLHF framework, recent RM studies have grown rapidly.
The original goal of RM is to provide a scalar score to a model response and indicate the quality in equation 2, especially helpfulness and harmlessness. Due to the trade-off in quality aspects (Touvron et al., 2023a; Bai et al., 2022b), it can be challenging for a single RM to perform well in all aspects. Our work related to previous works handling multiple rewards and potential disagreement in preferences. For instance, Llama-2 (Touvron et al., 2023a) utilizes two separate RMs, one optimized for helpfulness and another for harmlessness. They mitigate the magnitude bias of the reward scalar with a margin loss, which provides a large margin for pairs with distinct responses, and a smaller one for those with similar responses. Multiple RMs can be utilized as majority voting or averaging (Jaques et al., 2020; Jang et al., 2023) in the PPO (Schulman et al., 2017). Wang et al. (2023a) introduces a Bayesian-based approach called d-PM to align language model with human preferences with disagreement. Cheng et al. (2023a) proposes to train a customized RM from the general RM to avoid disagreement from different preference domains. Furthermore, our theoretical intuition follows recent work DPO (Rafailov et al., 2023) and SLiC-HF (Zhao et al., 2023b) for preference alignment, which explores more straightforward methods to align language models with human preferences. Beyond the methodology, they have shown the RLHF framework is working as likelihood calibration tasks (Deng et al., 2020; Wang et al., 2023a; Azar et al., 2023), which proves that the reward values provided by the RM are also important.
6 Conclusion
This research paper delves into the problem of reward modeling within the context of mixed diverse preferences datasets. Our experimental findings indicate that RMs trained on different preference datasets exhibit substantial diversity in their reward distributions. This diversity stems from the inherent characteristics of the datasets, posing a challenge for the RM to establish a consensus or accurately represent a shared preference. To tackle this challenge, we propose a straightforward yet highly effective training policy called MORE. The MORE policy is specifically designed to enhance the reward modeling process on diverse preference datasets, to improve the model’s ability to capture and represent the underlying preferences accurately. Through experiments conducted on five diverse datasets, we demonstrate the practical challenges associated with reward modeling. Furthermore, our proposed method outperforms conventional training policies across three different evaluation metrics, highlighting its effectiveness and superiority.
Appendix A Appendix
A.1 Decomposition of ranking loss
We present the detailed decomposing of vanilla ranking loss:
$\nabla_{\theta}\mathcal{L}_{\text{rank}}(\theta;\mathcal{B})$ $=\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\nabla_{\theta}r_{\theta}(x^{(b)},y_{w}^{(b)})-\nabla_{\theta}r_{\theta}(x^{(b)},y_{l}^{(b)})\right]$
$=\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\nabla_{\theta}(r_{\theta}^{*}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)}))\right]$
$\quad+\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\nabla_{\theta}\tilde{r}_{\theta_{i}}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)}))\right]$
$=\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\nabla_{\theta}(r_{\theta}^{*}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)})\right]$
$\quad+\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\sum_{i=1}^{K}\frac{1}{K}\nabla_{\theta}\tilde{r}_{\theta_{i}}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)}))\right]$
$=\sum_{b=1}^{B}\left[-\sigma\left(r_{\theta}(x^{(b)},y_{l}^{(b)})-r_{\theta}(x^{(b)},y_{w}^{(b)})\right)\right]\cdot\left[\nabla_{\theta}r_{\theta}^{*}(x^{(b)},y_{w}^{(b)}-y_{l}^{(b)})\right]$
$\quad+K\sum_{i=1}^{K}\frac{1}{K}\sum_{(x,y_{w},y_{l})\in\mathcal{B}_{i}}\left[-\sigma\left(r_{\theta}(x,y_{l})-r_{\theta}(x,y_{w})\right)\right]\cdot\left[\nabla_{\theta}\tilde{r}_{\theta}(x,y_{w}-y_{l}))\right],$
where we let $r_{\theta}(x,y_{w}-y_{l})=r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$. In contrast, the gradient $\nabla_{\theta}\mathcal{L}_{\text{MORE}}(\theta;\mathcal{B})$ replaces the coefficients $\frac{1}{K}$ with adjustable variable $\lambda$. Therefore, the vanilla ranking loss is a special care of MORE loss.
A.2 Experiments on Two Diverse Preference Datasets
In our preliminary experiments, we train RMs on two diverse preference datasets, including Helpful-Harmless and Webgpt-Oasst1. We use the full size of datasets from the original datasets. In details, For Helpful, Harmless, Webgpt and Oasst1, we have 43,774, 42,537, 17.106 and 18,165 training samples correspondingly. The number of training samples and test samples of each group is balanced. We present the results in the tables below.
Table 3: Performance on Webgpt-Oasst1 datasets.
Table 4: Performance on Helpful-Harmless datasets.
Webgpt OASST1 Avg. Acc ECE Raw 51.15 49.84 50.50 0.234 Webgpt 67.37 64.00 65.68 0.103 OASST1 66.31 66.61 66.46 0.102 Hybrid 66.43 65.88 66.15 0.101 Sequential 65.51 67.53 66.52 0.113 MORE 67.76 67.22 67.49 0.130
Helpful Harmless Avg. Acc ECE Raw 49.78 47.18 48.48 0.244 Helpful 66.75 35.89 51.32 0.300 Harmless 35.67 67.21 51.44 0.312 Hybrid 63.77 63.45 63.61 0.032 Sequential 64.15 63.40 63.75 0.028 MORE 64.58 63.14 63.86 0.051
Table 4: Performance on Helpful-Harmless datasets.
The results show that MORE implements comparable results with baseline methods in reward accuracy. However, the MORE is not the best on the ECE metric. This is because the number of diverse datasets is small (K=2), resulting in less diverse preference information. Hence, the reward drift mitigation task is solved inefficiently, making MORE degenerate to the vanilla method with rank loss as we demonstrated in Section A.1. Furthermore, we have achieved better results as shown in the main paper with $K=5$, indicating that MORE is efficient in handling diverse preferences in reward modeling.
A.3 Diversity of Reward Values
We provide missing experimental results as supplements of Figure 1.